OpenStudy (anonymous):

If the mass of a sample of radioactive material is given by m=800*2^(-0.04t)... grams, time in years a) What is the initial mass? b) After how long time is the mass 50g

OpenStudy (anonymous):

a: Initial mass should be 800 grams.

OpenStudy (anonymous):

Yes, I realise this. More interested in b), thanks (:

OpenStudy (anonymous):

a) 800 b) 100 years

OpenStudy (anonymous):

Solution?

OpenStudy (anonymous):

solution is to solve \[800\times2^{-0.04t}=50\] for \(t\)

OpenStudy (anonymous):

step 1) is to divide both sides by 800 and get \[2^{-0.04t}=\frac{1}{16}\]

OpenStudy (anonymous):

step 2) is to recognize that \(\frac{1}{16}\) is an integer power of 2, namely -4 because \(2^{-4}=\frac{1}{16}\) and final step is to set \[-0.04t=-4\] and solve for \(t\)